# Rapidly Converging Time-Discounted Ergodicity on Graphs for Active Inspection of Confined Spaces

### v1

@@ VII-A Graph World Experiment
[…]
@@ VII-A1 Methodology
The target distribution $\bar{\rho}$ and initial distribution $\rho_{0}$ are generated by uniformly sampling the weight for each region from $0-1$ and normalizing them so that they add up to $1$. The expected ergodicity deviation at each time step, $\|\mathbb{E}[\hat{\rho}_{k}]-\bar{\rho}\|_{2}$, is calculated both for FMMC and REMC, where
@@
[…]
@@ VII-A2 Results
Fig. 8 shows the expected ergodicity deviations in 10 time steps for FMMC and REMC separately for 1000 trials. The result shows that on average the REMC method attains ergodicity faster than FMMC. Additionally, it has a substantially lower worst case deviation (shown by the top blue whiskers) than FMMC.
@@
Fig. 9 shows the pairwise difference (FMMC - REMC) between the two methods for the same trials. The result shows that FMMC is consistently less ergodic, with the first quartiles being larger than 0 at at all the time steps.
@@ VII-B Grid World Experiment
The following experiment shows the comparison of region-based ergodicity between HEMaP and a continuous space ergodic planner, spectral multi-scale coverage (SMC), in the occupancy grid of the ballast tank shown in Fig. 2.
@@ VII-B1 Methodology
The occupancy grid map is created by projecting the ballast tank CAD model onto the ground, with a lower and upper height limit of $0.05$ m and $0.2$ m, respectively. Additionally, the occupied grids are inflated with the radius of a TurtleBot Burger. The robot is modeled as a point mass moving at a constant speed of 0.2 m/s. The target distribution is uniform for each region, but varies between the regions, as shown in Fig. 10. To account for collision avoidance, the grid cells belonging to obstacles are considered an extra region, with a target visitation frequency of 0. For SMC, 20 Fourier coefficients are used with a time step of 0.01 s. For the hierarchical method, the waypoints are selected uniformly randomly in the target region indicated by the graph planner. A* is chosen as the path planning algorithm to navigate between the waypoints. The robot is allowed to run for a total of 1,200 s and a total of 30 trials are performed for each method.
@@
Figure 10: Example ergodic trajectories for HEMaP and SMC, with the darker blue shade indicating higher target frequency. The green circles represent the initial location; the black dots in (a) represent the inspection waypoints and dashed gray line shows intermittent trajectory; the gray line in (b) shows continuous trajectory. SMC yields more complex trajectories that enter the obstacle region even if the desired distribution is set to 0.
VII-B2 Results
@@
Fig. 10 shows example trajectories for both SMC and HEMaP. As shown in Fig. 10(a), the combination of graph planner and A* is able to generate a trajectory that is completely contained in the workspace. In contrast, since SMC only considers a finite number of Fourier coefficients, it cannot always avoid the obstacles. Even worse, the Gibbs phenomenon causes the robot to actively run into obstacles, as seen by the long arcs stretching outside the workspace. In a similar vein, the SMC does not have any knowledge of the topology of the space, which allows the robot to pass through obstacles and reach different regions directly. Additionally, the path generated by A* is more straightforward and easier for a ground robot to follow than the complex trajectory generated by SMC.
@@
Figure 11: Comparison of region based ergodicity between SMC and HEMaP in the tank in Fig. 3 with the robot running at 0.2 m/s for 1,200 seconds. HEMaP has better region-based ergodicity on average, but with higher variance.
Figure 12: Pair-wise error difference from each trial calculated from the data from Fig. 11. For the same initial condition and target distribution, HEMaP consistently has less error on average compared to SMC, with the 3rd quartile above 00 after 200 seconds.
Fig. 11 shows region-wise ergodic deviation at each time step and Fig. 12 shows the pairwise difference (SMC - HEMaP) for each trial on the same dataset. For SMC, since it is a continuous space method, the relative frequency is calculated based on the current location (region) of the robot, with obstacles treated as an extra region. For HEMaP, since it is a discrete space planner, the relative frequency is calculated for each waypoint and the ergodicity deviation is held constant from the previous waypoint while the robot is in transition. The result shows that HEMaP is able to reduce the ergodicity deviation faster than SMC on average, even when SMC is allowed to ignore the topology of the explored space. This can be explained by the fact that SMC is designed to handle rectangular domains. SMC partitions the space into regular grid-like regions at varying scales using a Fourier basis. However, the tank regions are irregular and non-rectangular with sharp changes in the desired distribution values at the region boundaries, which is hard to capture using a Fourier series. [^1] If a more advanced continuous space ergodic planner is used with a careful consideration of collision avoidance, the error from the obstacle region should be greatly reduced. However, the problem of choosing finite Fourier coefficients to approximate the region-based target distribution will persist. HEMaP, however, has a higher variance across the trials due to its stochastic nature, while SMC is relatively consistent.
@@ VII-C Gazebo Simulation
The following experiment compares the performance of two different graph traversal methods against REMC ergodic traversal, in the context of FOD detection in the ballast tank of Fig. 2. Specifically, we consider random walk, where the planner samples from all the edges in each region uniformly; and a greedy method, where the planner chooses the edge that leads to the neighboring region with the highest estimated information measure. Additionally, the waypoint placements within the regions are changed accordingly, with random walk choosing the sampled waypoints randomly; and the greedy method choosing the sampled waypoints with the maximum estimated information.
@@ VII-C1 Methodology
Simulation is performed in Gazebo using the ballast tank model in Fig. 2. A rusty texture is applied to the entire tank structure and all the light sources are removed. The robot is a TurtleBot3 Burger with the original camera replaced by a RealSense D435. AprilTags are scattered around the tank as partially known features and the robot wheel encoders provide odometry information for the hierarchical SLAM framework described in Section V-A. A representative image of the simulated environment is shown in Fig. 13. The same occupancy grid map from Section VII-B is used as the static global cost map for the ROS navigation planner and a LiDAR is used to generate the local cost map for the local planner. The number of points in the reference point cloud, $\mathcal{P}_{\text{ref}}$, is empirically selected as $100,000$ to provide sufficient density for the detection of FODs, given the size of the tank and the FODs. The buffer zone threshold, $\epsilon$, in Fig. 6 is selected to be $3.5$ cm. The smoothing factor, $\delta$, in Algorithm 1 is selected to be $0.001$.
@@ VII-C2 Results
Table I shows the results for the FODs detected in the photos by the three inspection methods over 15 trials. A total of 76 FODs are present, of which the ergodic method only misses 12 FODs. However, the random and greedy methods miss a much larger number of FODs. This occurs due to an inherent bias toward visiting highly connected regions and high-information regions by the random and greedy methods, respectively. As a result, some of the other tank regions are not inspected adequately when the total number of time steps is limited. The visitation frequency for each region is discussed in more detail in our subsequent analysis. The ergodic and random methods share a similar number of false positive photos, while the greedy has the lowest number. This happens because both HEMaP and random methods distribute the inspection points, whereas, the greedy method focuses on refining several high information points.
@@
Furthermore, 15 trials with the same setup are performed for 15 and 25 steps, and the detection rates are shown in Fig. 16. As expected, all the methods have a decreased detection rate with a smaller number of steps. Nevertheless, even with a small number of steps, ergodic traversal is more effective in detecting FODs than the other two methods. This experiment showcases the adaptability of REMC to arbitrary target distributions and how it is combined effectively with the anomaly detector in the HEMaP framework. The random and greedy exploration methods are highly dependent on the graph structure, where the visitation frequencies are determined by the connectivity of the graph, causing the highly connected regions to be visited more frequently. Additionally, the convergence rates of the time averages to the corresponding target distributions are not optimized for the random and greedy methods. Therefore, the short-term visitation frequency is biased toward the regions closer to the starting region, while the far away regions are less likely to be visited.
@@ VII-D1 Methodology
The hierarchical SLAM is run once in teleoperation mode to map the AprilTags locations before the autonomous inspection trials. The depth point clouds are also collected and assembled into a global map to generate the reference CAD model. Similar to Section VII-C, the occupancy grid map is generated by projecting the reference CAD model within a height range of $0.05$ m to $0.2$ m. Similarly, the occupancy grid map is used for the global planner and the LiDAR is used for the local planner. The number of points in the reference point cloud, $\mathcal{P}_{\text{ref}}$, is empirically selected as $100,000$. The buffer zone threshold, $\epsilon$ and the smoothing factor, $\delta$, are chosen to be $1$ cm and $0.001$, respectively. A set of six common hand tools is chosen for the FODs, specifically: 1) power drill $\times$ 2; 2) masking tape; 3) sander; 4) crimper; 5) sander. A total of 5 trials are performed with 35 steps each. $3$ to $5$ FODs were sampled for each trial with a total of $21$ FODs for the five trials. Each FOD is placed in the tank by uniformly sampling from the edges of the occupancy map. If the FODs are hidden behind a structure or blocking the doorways, they are moved to the nearest available location.
@@ VIII Discussion
The high variance in performance arising from the stochasticity of the Markov chain can be reduced through Monte Carlo tree search (MCTS) [13]. In MCTS, multiple sequences with a predefined time horizon are sampled from the Markov chain, and the sequence with the highest increase in ergodicity is selected instead of sampling the next state at every step. Here, we use the single-step sampling option to investigate the fundamental capability of an ergodic Markov chain and devise an exploration framework that can be extended to communication-limited multi-robot systems. This extended framework can be potentially applied to a wider array of exploration problems beyond inspection, such as wildlife monitoring in a biome and lost object retrieval in a building environment.
@@
It is also possible to consider an ergodic Markov chain in the state space instead of the physical space. The state space can be either graph-based or topological, with inherently stochastic dynamics analogous to a Markov decision process. A classical strategy for state space exploration in reinforcement learning (RL) is to uniformly sample an available action [55, p.32]. This is analogous to the random method in Section VII-C, where the visitation frequency is highly dependent on the connectivity of the space. Therefore, we are interested in investigating whether an on-policy RL method can be developed using an ergodic Markov chain-based stochastic policy.
@@ IX Conclusions
This paper addresses the problem of efficient active robotic inspection of confined spaces using anomaly detection as a case study. First, an ergodic exploration method, named rapidly ergodic Markov chain (REMC), is introduced for a graph-based discretization of the confined space. REMC is shown to have an optimal convergence rate to time-discounted ergodicity on any reversible graph, implying it prioritizes early visitations of information-rich graph nodes (regions). REMC also provides an upper bound on the convergence rate to time-discounted ergodicity for any strongly connected graph. Therefore, REMC achieves more meaningful ergodicity in fewer time steps than the classical fastest mixing Markov chain method, which is optimized for the convergence rate to the target distribution. Second, a hierarchical ergodic Markov planner (HEMaP) is developed that uses REMC for region graph traversal, a waypoint planner to select the inspection points inside the regions, and a standard local planner for collision-free waypoint navigation. HEMaP is empirically shown to have better convergence to region-based ergodicity than spectral multi-scale coverage (SMC), a widely used continuous space ergodic control method. Finally, a SLAM-driven Bayesian anomaly detection method is incorporated within HEMaP for robust localization of foreign object debris (FODs) in the confined space. Both simulation and physical experiments in a ballast tank using a ground robot show that HEMaP leads to a significantly higher FOD detection rate compared to greedy and random inspection methods.

### v2

@@ VII-A Graph World Experiment
The following experiment compares the performance between the ergodicity of the Markov chain generated by FMMC and REMC from Theorem 3 using the region graph in Fig. 3.
@@ VII-A1 Methodology
The target distribution $\bar{\rho}$ and initial distribution $\rho_{0}$ are generated by uniformly sampling the weight for each region from 010-1 and normalizing them so that they add up to $1$. The expected ergodicity deviation at each time step, $\|\mathbb{E}[\hat{\rho}_{k}]-\bar{\rho}\|_{2}$, is calculated both for FMMC and REMC, where
@@
Figure 8: Ergodicity comparison between fastest mixing Markov chain (FMMC) and REMC from Theorem 3 tested on the region graph from Fig. 3, showing that REMC has a lower ergodicity error with a small number of steps on average.
@@ VII-A2 Results
Fig. 8 shows the expected ergodicity deviations in 10 time steps for FMMC and REMC separately for 1,000 trials. The result shows that on average the REMC method attains ergodicity faster than FMMC. Additionally, it has a substantially lower worst case deviation (shown by the top blue whiskers) than FMMC. Fig. 9 shows the pairwise difference (FMMC - REMC) between the two methods for the same trials. The result shows that FMMC is consistently less ergodic, with the first quartiles larger than 0 at all the time steps.
@@ VII-B Grid World Experiment
The following experiments shows the comparison of region-based ergodicity between HEMaP and three other continuous space ergodic planners: spectral multiscale coverage (SMC) [9]; heat equation driven area coverage (HEDAC) [20]; and graph-based ergodic search in cluttered environments (GESCE) [31], in the occupancy grid of the ballast tank shown in Fig. 2.
@@ VII-B1 Methodology
The occupancy grid map is created by projecting the ballast tank CAD model onto the ground, with a lower and upper height limit of $0.05$ m and $0.2$ m, respectively. Additionally, the occupied grids are inflated with the radius of a TurtleBot Burger. The robot is modeled as a point mass moving at a constant speed of 0.2 m/s. Specifically, for SMC, the robot follows a single integrator model
$\begin{bmatrix}x_{t+1}\\ y_{t+1}\end{bmatrix}=\begin{bmatrix}x_{t}\\ y_{t}\end{bmatrix}+\Delta t\begin{bmatrix}v_{x}\\ v_{y}\end{bmatrix}$ (78)
with $\sqrt{v_{x}^{2}+v_{y}^{2}}=0.2$. For HEMaP, the robot is allowed to move to any of the 8 neighboring grid cells centered at the current grid location within the time interval $\Delta t$. For HEDAC, the robot follows second-order dynamics according to this GitLab implementation [^2] https://gitlab.idiap.ch/rli/robotics-codes-from-scratch/-/blob/master/python/ergodic_control_HEDAC_2D.py?ref_type=heads. For GESCE, the robot travels directly according to a probabilistic road map (PRM) planner, i.e., it follows a straight line from one collision-free node to the next without any low-level dynamical constraints.
The target distribution is uniform for each region, but varies among the regions, as shown in Fig. 10. To account for collision avoidance, grid cells belonging to obstacles are considered an extra region, with a target visitation frequency of 0. For SMC, 20 Fourier coefficients are used with a time step of 0.01 s. For HEMaP, the waypoints are selected uniformly randomly in the target region indicated by the graph planner. A* is chosen as the path planning algorithm to navigate between the waypoints. For GESCE, 200 nodes are uniformly randomly sampled from the free space for the PRM with any pairs of nodes within 0.6 m connected by an edge. The cost function is identical to the SMC implementation. A total of 500 iterations are allowed during the search cycle to maintain a reasonable runtime. The robot is allowed to run for a total of 1,200 s (except for GESCE), and a total of 30 trials are performed for each method.
VII-B2 Results
Figure 10: Example trajectories generated by HEMaP with A* as the low level planner and uniformly random waypoint placement. This shows that ergodic planning can adapt to various spatial distributions $\bar{\rho}$, where the number of sampled points (shown as black circles) is proportional to the desired region visitation frequency (shown as shade of blue).
@@ VII-B2 Results
Figure 11: Example ergodic trajectories comparison for HEMaP and SMC, with the darker blue shade indicating higher target frequency. The green circles represent the initial location; the black dots in (a) represent the inspection waypoints and dashed gray line shows intermittent trajectory; the gray line in (b) shows continuous trajectory. SMC yields more complex trajectories that enter the obstacle region even if the desired distribution is set to 0.
@@
A few representative robot inspection trajectories are shown in Fig. 10, which demonstrate that HEMaP works well regardless of the target distribution $\bar{\rho}$ that varies according to the region-level anomaly entropy. To provide a contrast with the continuous space ergodic planner trajectories, Fig. 11 shows comparisons between HEMaP and SMC for the same choice of $\bar{\rho}$. While HEMaP trajectory is completely contained in the workspace, SMC cannot always avoid the obstacles as it considers a finite number of Fourier coefficients. The Gibbs phenomenon also causes the robot to actively run into obstacles, as seen by the long arcs stretching outside the workspace. In a similar vein, SMC does not have any knowledge of the topology of the space, which allows the robot to pass through obstacles and reach different regions directly. Additionally, the path generated by the A* planner in HEMaP is more straightforward for a ground robot to follow than the complex trajectory generated by SMC.
Fig. 12 shows region-wise ergodic deviations at each time step for all the ergodic planners averaged over the complete set of trials. Fig. 13 plots the statistics on the pairwise differences between HEMaP and the continuous space planners on the same set of trials. For SMC and HEDAC, the relative frequency is calculated based on the current location (region) of the robot, with obstacles treated as an extra region. For HEMaP, the relative frequency is calculated for each waypoint and the ergodicity deviation is held constant from the previous waypoint while the robot is in transition. Similarly, the relative frequency for GESCE is calculated for each node in the PRM graph and held constant during the traversal between the nodes.
@@
Figure 12: Comparison of region-based ergodicity between HEMaP and the various continuous space ergodic planners in the ballast tank (shown in Fig. 3) with the robot running at 0.2 m/s for 1,200 seconds. HEMaP has better region-based ergodicity on average.
(a) SMC
(b) HEDAC
(c) GESCE
Figure 13: Statistics on pair-wise differences in ergodicity deviations (errors) between HEMaP and the three continuous space planners over the same ser of trials used in Fig. 12. (a) shows that for the same initial condition and target distribution, HEMaP consistently has less error on average compared to SMC, with the 3rd quartile above 0 after 200 seconds. (b) shows a similar trend for HEDAC, where HEDAC has a higher initial error but lower asymptotic error because of its collision avoidance capability. (c) shows GESCE has an almost identical performance to HEMaP in terms of the 1st quartile values (worse with respect to the median and 3rd quartile values though), but is unable to run for more than 200 seconds.
The results show that HEMaP is able to reduce the region-based ergodicity deviation faster than all the other planners on average. This happens even though SMC does not consider the topology of the explored space, and can be explained by the fact that SMC is designed to handle rectangular domains that are partitioned into regular grid-like regions at varying scales using a Fourier basis. However, the ballast tank regions are irregular and non-rectangular with sharp changes in the desired distribution values at the region boundaries, which are hard to capture using a Fourier series. HEDAC, on the other hand, has a higher error than SMC initially due to obstacle avoidance. However, it is capable of approaching the performance of HEMaP at around 800 s, since it has significantly less collisions compared to SMC [^3] Minor collisions, similar to Fig. 4 in [20], can still be observed. This is likely due to discretized second-order dynamics, where the potential field does not produce sufficient acceleration for collision avoidance.. GESCE has similar initial performance as SMC while maintaining obstacle avoidance, and improves substantially if the search is allowed to run longer. The main challenge is that the trajectory length cannot be specified, due to which none of the trials exceed the 200 second mark, while all the other methods can run indefinitely.
In general, the main advantage of HEMaP is that it discards fine-grained information in pursuit of rapid convergence to the target visitation distribution of inspection regions. Therefore, it takes the shortest A* path between the regions, which enhances ergodicity at the region level. This is in contrast to all the continuous space methods that try to provide intra-regional coverage, which affects region-level ergodicity deviation and leads to fractal-like trajectories. If the fine-grained information are crucial and the robot is able to execute complex trajectories, then one can consider a continuous space method. Alternatively, since HEMaP is modular and hierarchical, we envision the possibility of replacing the A* planner with any compatible continuous space method, which is further discussed in Section VIII.
@@ VII-C Gazebo Simulation
The following experiment compares the performance of two different graph traversal methods against HEMaP, in the context of FOD detection in the ballast tank of Fig. 2. Specifically, we consider random walk, where the planner samples from all the edges in each region uniformly; and a greedy method, where the planner chooses the edge that leads to the neighboring region with the highest estimated information measure. Additionally, the waypoint placements within the regions are changed accordingly, with random walk choosing the sampled waypoints randomly; and the greedy method choosing the sampled waypoints with the maximum estimated information.
@@ VII-C1 Methodology
Simulation is performed in Gazebo using the ballast tank model in Fig. 2. A rusty texture is applied to the entire tank structure and all the light sources are removed. The robot is a TurtleBot3 Burger with the original camera replaced by a RealSense D435. AprilTags are scattered around the tank as partially known features and the robot wheel encoders provide odometry information for the hierarchical SLAM framework described in Section V-A. A representative image of the simulated environment is shown in Fig. 14. The same occupancy grid map from Section VII-B is used as the static global cost map for the ROS navigation planner and a LiDAR is used to generate the local cost map for the local planner. The number of points in the reference point cloud, $\mathcal{P}_{\text{ref}}$, is empirically selected as $100,000$ to provide sufficient density for the detection of FODs, given the size of the tank and the FODs. The buffer zone threshold, $d_{\text{buffer}}$, in Fig. 7 is selected to be $3.5$ cm. The smoothing factor, $\delta$, in Algorithm 1 is selected to be $0.001$.
@@ VII-C2 Results
Table I shows the results for the FODs detected in the photos by the three inspection methods over 15 trials. A total of 76 FODs are present, of which HEMaP only misses 12 FODs. However, the random and greedy methods miss a much larger number of FODs. This occurs due to an inherent bias toward visiting highly connected regions and high-information regions by the random and greedy methods, respectively. As a result, some of the other tank regions are not inspected adequately when the total number of time steps is limited. The visitation frequency for each region is discussed in more detail in our subsequent analysis. HEMaP and random methods share a similar number of false positive photos, while the greedy has the lowest number. This happens because both HEMaP and random methods distribute the inspection points, whereas, the greedy method focuses on refining several high information points.
@@
Furthermore, 15 trials with the same setup are performed for 15 and 25 steps, and the detection rates are shown in Fig. 17. As expected, all the methods have a decreased detection rate with a smaller number of steps. Nevertheless, even with a small number of steps, HEMaP is more effective in detecting FODs than the other two methods. This experiment showcases the adaptability of REMC to arbitrary target distributions and how it is combined effectively with the anomaly detector in the HEMaP framework. The random and greedy exploration methods are highly dependent on the graph structure, where the visitation frequencies are determined by the connectivity of the graph, causing the highly connected regions to be visited more frequently. Additionally, the convergence rates of the time averages to the corresponding target distributions are not optimized for the random and greedy methods. Therefore, the short-term visitation frequency is biased toward the regions closer to the starting region, while the far away regions are less likely to be visited.
@@ VII-D1 Methodology
The hierarchical SLAM is run once in teleoperation mode to map the AprilTags locations before the autonomous inspection trials. The local point clouds are also collected and assembled into a global map to generate the reference CAD model. Similar to Section VII-C, the occupancy grid map is generated by projecting the reference CAD model within a height range of $0.05$ m to $0.2$ m. Similarly, the occupancy grid map is used for the global planner and the LiDAR is used for the local planner. The number of points in the reference point cloud, $\mathcal{P}_{\text{ref}}$, is empirically selected as $100,000$. The buffer zone threshold, $d_{\text{buffer}}$ and the smoothing factor, $\delta$, are chosen to be $1$ cm and $0.001$, respectively. A set of six common hand tools is chosen for the FODs, specifically: 1) power drill $\times$ 2; 2) masking tape; 3) sander; 4) crimper; 5) sander. A total of 5 trials are performed with 35 steps each. $3$ to $5$ FODs were sampled for each trial with a total of $21$ FODs for the five trials. Each FOD is placed in the tank by uniformly sampling from the edges of the occupancy map. If the FODs are hidden behind a structure or blocking the doorways, they are moved to the nearest available location.
@@ VIII Discussion
The high variance in performance arising from the stochasticity of the Markov chain can be reduced through Monte Carlo tree search (MCTS) [13], although as shown in Appendix C, the variance of the time average of REMC decreases rapidly with the number of steps. In MCTS, multiple sequences with a predefined time horizon are sampled from the Markov chain, and the sequence with the highest increase in ergodicity is selected instead of sampling the next state at every step. Here, we use the single-step sampling option to investigate the fundamental capability of an ergodic Markov chain and devise an exploration framework that can be extended to communication-limited multi-robot systems. This extended framework can be potentially applied to a wider array of exploration problems beyond inspection, such as wildlife monitoring in a biome and lost object retrieval in a building environment.
@@
It is also possible to consider an ergodic Markov chain in the state space instead of the physical space. The state space can be either graph-based or topological, with inherently stochastic dynamics analogous to a Markov decision process. A classical strategy for state space exploration in reinforcement learning (RL) is to uniformly sample an available action [14, p.32]. This is analogous to the random method in Section VII-C, where the visitation frequency is highly dependent on the connectivity of the space. Therefore, we are interested in investigating whether an on-policy RL method can be developed using an ergodic Markov chain-based stochastic policy.
@@ IX Conclusions
This paper addresses the problem of efficient active robotic inspection of confined spaces using anomaly detection as a case study. First, an ergodic exploration method, named rapidly ergodic Markov chain (REMC), is introduced for a graph-based discretization of the confined space. REMC is shown to have an optimal convergence rate to time-discounted ergodicity on any reversible graph, implying it prioritizes early visitations of information-rich graph nodes (regions). REMC also provides an upper bound on the convergence rate to time-discounted ergodicity for any strongly connected graph. Therefore, REMC achieves more meaningful ergodicity in fewer time steps than the classical fastest mixing Markov chain method, which is optimized for the convergence rate to the target distribution. Second, a hierarchical ergodic Markov planner (HEMaP) is developed that uses REMC for region graph traversal, a waypoint planner to select the inspection points inside the regions, and a standard local planner for collision-free waypoint navigation. HEMaP is empirically shown to have better convergence to region-based ergodicity than a widely-used and two more-recent, continuous space ergodic control methods (SMC, HEDAC and GESCE, respectively).Finally, a SLAM-driven Bayesian anomaly detection method is incorporated within HEMaP for robust localization of foreign object debris (FODs) in the confined space. Both simulation and physical experiments in a ballast tank using a ground robot show that HEMaP leads to a significantly higher FOD detection rate compared to greedy and random inspection methods.